OpenStudy (anonymous):

Assume the y varies inversely with x and y = -3 when x = 4. Write the indirect variation equation for the relationship. a. xy = -3/4 b. xy = 4/-3 c. xy = -12

OpenStudy (gorv):

@AlexusFerilli c is the ans need explanation???

OpenStudy (anonymous):

Thank you and yes please.

OpenStudy (gorv):

let y=a/x so y=-3 then x=4 implies xy=a -3*4=a a=-12 y=-12/x xy=-12 plz give best response if u got it